OpenStudy (anonymous):

Which equations in point-slope form are equations of the line that pass through the points (4, 5) and (−3, −1)? Choose exactly two answers that are correct. A. y - 3 = 7/6(x - 1) B. y - 5 = 6/7(x - 4) C. y + 1 = 6/7(x + 3) D. y - 4 = 7/6(x - 5)

OpenStudy (anonymous):

@GabeBae Give me a hand?

OpenStudy (anonymous):

Use this formula\[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

First,find the gradient/slope. The formula for slope is\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]\[m=\frac{ -1-5 }{ -3-4 }\]\[m=\frac{ -6 }{ -7 }\]\[m=\frac{ 6 }{ 7 }\]

OpenStudy (anonymous):

Then use both of the points \[(4,5)and(-3,-1)\]to find two equations

OpenStudy (anonymous):

First equation\[First~point,(4,5)=(x_1,y_1)\]\[m=\frac{ 6 }{ 7 }\]Substitute the value into this formula equation\[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

Second equation\[Seccond~point,(-3,-1)=(x_1,y_1)\]\[m=\frac{ 6 }{ 7 }\]Substitute the value into this formula equation\[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

@sonic400539